suppose a firm has a cobb-douglas weekly production function q=f(l,k)=25l^.5k^.5, where l is the number of workers and k is units of capital.mrtslk is k/l. the wage rate is $900 per week, and a unit of capital costs $400 per week. what is the least cost input combination for producing 675 units of output?
Answer #1

Least cost combination is determined where MRTS = wage rate/capital cost = 900/400 = 2.25

MRTS = MPL/MPK = 25(0.5)0.5-10.5 25(0.5)050.5- 1-0.510.5 10.5-0.5 A:0.5+0.5 Te 0.5+0.5

So, MRTS = 2.25 gives,
k/l = 2.25
So, k = 2.25l

q = 675 = 25l^.5k^.5 = 25l.5(2.25l).5 = 25l.5(1.5)(l).5 = 37.5l.5+.5 = 37.5l
So, l = 675/37.5 = 18
k = 2.25l = 2.25*(18) = 40.5

So, l = 18; k = 40.5